# Uncertainty in Language Models: Assessment through Rank-Calibration

### v1

@@ Appendix B Common Uncertainty/Confidence Measures for LMs
    A natural extension accounts for the variable length of responses by applying length normalization. Supposing the number of tokens of the response $\widehat{\boldsymbol{y}}$ is ${\rm len}(\widehat{\boldsymbol{y}})$, the length-normalized NLL is defined as
@@ E.2 Datasets
TriviaQA is a challenging reading comprehension dataset, containing question-answer pairs whose answers can be found on Wikipedia and the web. Similar to previous works, we use TriviaQA as an open-domain QA benchmark. Natural Questions is a question-answering dataset containing questions issued to the Google search engine. We use Natural Questions as an open-domain QA benchmark. SQuAD-1 is a reading comprehension dataset containing questions posed by crowdworkers based on Wikipedia articles. We include SQuAD-1 as a reading comprehension benchmark, where the annotated contexts are provided in the prompt. Meadow is created by research groups working on COVID-19 problems. We use this dataset for open-ended generation, where the LM is expected to provide a title for a paper given the abstract of the paper. The correctness is justified by comparing the generated title to the true title.
@@ E.3 Correctness Functions
The Metric for Evaluation of Translation with Explicit Ordering (METEOR) score has also been originally designed to evaluate machine translation and text summarization. Different from the Rouge score, the METEOR score considers the accuracy and fluency of the generation, as well as word order. Calculation of METEOR score can be found in Banerjee and Lavie (2005).
@@ E.4 Inconsistency due to Correctness Thresholding
We provide more evidence to show the inconsistency of AUARC and AUPRC metrics caused by ad hoc correctness thresholding. The plots are in Fig 5,  6, 7, 8, and 9.
Refer to captionRefer to caption
@@ F.1 Qualitative Illustration
Refer to captionRefer to caption
(a) Meadow
Refer to captionRefer to caption
(b) NQ-Open
Refer to captionRefer to caption
(c) Squad
Refer to captionRefer to caption
(d) TrviaQA
@@ F.2 Recalibration with Histogram Binning
where $\mathcal{I}_{b,{\rm cal}}\triangleq\{i:1\leq i\leq n_{\rm cal},u_{i}\in{\rm bin}_{b}\}$. We re-calibrate the measure $U$, defining $U_{\rm cal}$ via $U_{\rm cal}(\boldsymbol{x};\widehat{\boldsymbol{y}})={\rm crc}_{b,{\rm cal}}$ for any $U(\boldsymbol{x};\widehat{\boldsymbol{y}})\in{\rm bin}_{b}$. We evaluate the performance of the calibrated measure on the test set. Table 4 and Fig. 10 and 11 list the RCE results of $U_{\rm SE}$ for GPT-3.5-turbo before and after calibration. We observe the calibrated measure is significantly better rank-calibrated, showing the effectiveness of this strategy. Here, we split the benchmark data equally into calibration and test sets.
Refer to captionRefer to caption
@@ F.3 Critical Difference Diagrams
dataset correctness temperature $U_{\rm SE}$ $U_{\rm SE,cal}$
meadow bert 1.0 0.177±0.027 0.083±0.016
meteor 1.0 0.132±0.018 0.066±0.015
rougeL 1.0 0.113±0.022 0.063±0.014
rouge1 1.0 0.113±0.018 0.061±0.012
nq-open bert 1.0 0.050±0.007 0.026±0.007
meteor 1.0 0.060±0.009 0.033±0.011
rougeL 1.0 0.052±0.008 0.030±0.010
rouge1 1.0 0.051±0.008 0.029±0.010
squad bert 1.0 0.113±0.013 0.050±0.013
meteor 1.0 0.086±0.014 0.046±0.010
rougeL 1.0 0.100±0.011 0.037±0.008
rouge1 1.0 0.103±0.011 0.039±0.007
triviaqa bert 0.5 0.052±0.009 0.030±0.010
1.0 0.052±0.012 0.027±0.008
1.5 0.081±0.009 0.029±0.007
meteor 0.5 0.234±0.019 0.058±0.015
1.0 0.209±0.012 0.047±0.014
1.5 0.176±0.015 0.036±0.012
rougeL 0.5 0.050±0.008 0.028±0.007
1.0 0.059±0.009 0.026±0.007
1.5 0.104±0.007 0.028±0.006
rouge1 0.5 0.050±0.008 0.028±0.006
1.0 0.060±0.009 0.027±0.006
1.5 0.105±0.008 0.028±0.008

### v2

@@ Appendix B Common Uncertainty/Confidence Measures for LMs
    A natural extension accounts for the variable length of responses by applying length normalization. Suppose that the number of tokens of the response $\widehat{\boldsymbol{y}}$ is ${\rm len}(\widehat{\boldsymbol{y}})$, the length-normalized NLL is defined as
@@ E.2 Datasets
TriviaQA is a challenging reading comprehension dataset, containing question-answer pairs whose answers can be found on Wikipedia and the web. Similar to previous works, we use TriviaQA as an open-domain QA benchmark. Natural Question is a question-answering dataset containing questions issued to the Google search engine. We use Natural Questions as an open-domain QA benchmark. SQuAD-1 is a reading comprehension dataset containing questions posed by crowdworkers based on Wikipedia articles. We include SQuAD-1 as a reading comprehension benchmark, where the annotated contexts are provided in the prompt. Meadow is created by research groups working on COVID-19 problems. We use this dataset for open-ended generation, where the LM is expected to provide a title for a paper given the abstract of the paper. The correctness is justified by comparing the generated title to the true title.
@@ E.3 Correctness Functions
The Metric for Evaluation of Translation with Explicit Ordering (METEOR) score has also been originally designed to evaluate machine translation and text summarization. Different from the Rouge score, the METEOR score considers the accuracy and fluency of the generation, as well as word order. The calculation of the METEOR score can be found in Banerjee and Lavie (2005).
@@ E.4 Inconsistency due to Correctness Thresholding
We provide more evidence to show the inconsistency of AUARC and AUPRC metrics caused by the ad hoc correctness thresholding. The plots are in Fig 5,  6, 7, 8, and 9.
@@ F.2 Recalibration with Histogram Binning
where $\mathcal{I}_{b,{\rm cal}}\triangleq\{i:1\leq i\leq n_{\rm cal},u_{i}\in{\rm bin}_{b}\}$. We re-calibrate the measure $U$, defining $U_{\rm cal}$ via $U_{\rm cal}(\boldsymbol{x};\widehat{\boldsymbol{y}})={\rm crc}_{b,{\rm cal}}$ for any $U(\boldsymbol{x};\widehat{\boldsymbol{y}})\in{\rm bin}_{b}$. We split the benchmark data equally into calibration and test sets and evaluate the performance of the calibrated measure on the test set. Table 3 and Fig. 10 and 11 list the RCE results of $U_{\rm SE}$ for GPT-3.5-turbo before and after calibration. We observe the calibrated measure is significantly better rank-calibrated, showing the effectiveness of this strategy.
(a) Meadow
(b) NQ-Open
(c) Squad
(d) TrviaQA
@@ F.4 Robustness Analysis
Influence of sample size.
We show that the empirical RCE is robust regarding the influence of sample size, which is crucial in scenarios where labeled data is hard to acquire. To this end, we conducted a new experiment using less data in the RCE computation, simulating scenarios where only a small amount of labeled data can is available. Specifically, we utilize 20%, 40%, 80%, 100% of the TriviaQA dataset in computing the empirical RCE values of uncertainty/confidence measures for the GPT-3.5 model with temperature 1.0. The RCE results under the Bert-similarity and RougeL correctness are in Table 6. The binning scheme is the same as the one used in the paper (i.e., 20 equal-mass bins). From the new experimental results, we observe that the RCE results are fairly stable, up to reasonable standard deviations (denoted by the subscript numbers), for moderately large datasets.
Proportion Correctness $U_{\rm Ecc}$ $U_{\rm Deg}$ $U_{\rm EigV}$ $U_{\rm NLL}$ $U_{\rm SE}$ $C_{\rm Verb}$
bert 20% 0.176±0.022 0.153±0.023 0.152±0.024 0.058±0.009 0.080±0.015 0.254±0.042
40% 0.171±0.020 0.151±0.021 0.154±0.020 0.048±0.010 0.083±0.013 0.211±0.045
80% 0.162±0.022 0.153±0.016 0.151±0.017 0.043±0.010 0.062±0.012 0.203±0.031
100% 0.152±0.025 0.129±0.020 0.133±0.020 0.039±0.007 0.052±0.012 0.182±0.025
rougeL 20% 0.178±0.020 0.153±0.024 0.153±0.023 0.061±0.010 0.098±0.016 0.238±0.035
40% 0.172±0.022 0.153±0.021 0.156±0.017 0.048±0.009 0.090±0.010 0.194±0.040
80% 0.156±0.020 0.145±0.017 0.146±0.017 0.042±0.009 0.073±0.013 0.190±0.030
100% 0.151±0.024 0.126±0.019 0.129±0.019 0.038±0.007 0.059±0.009 0.181±0.026
Table 6: RCE results for GPT-3.5-turbo (temperature 1.0) performing on the TriviaQA data with various dataset sizes under the Bert-similarity and RougeL correctness.